Cost of Capital - CAPM

The Inception, Inc wishes to calculate the cost of common stock equity. The analysts indicate that the risk-free is 5%, the prime rate is 8%, the firm's beta is 1.2 and the market rate is 12%.

Please show the work.
Answer #1

Cost of Equity = risk free rate+(market rate-risk free rate)*beta

= 5+(12-5)*1.2

= 5+ 7*1.2

= 5+8.4

= 13.4

Therefore cost of equity is 13.4%